OpenStudy (anonymous):

how do you find the root of this equation, 2x squared-x-15=0

hartnn (hartnn):

@sdahar hi:) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) do you know how to solve quadratic equations ?

hartnn (hartnn):

using factorization ?

OpenStudy (anonymous):

i really need help with slopes and how to find them in equations

hartnn (hartnn):

your original question needs to find roots, right ? can you factorize that quadratic by splitting the middle term ?

OpenStudy (campbell_st):

you can factorise it... or use the general quadratic formula. here is a fail safe method for factorising for a quadractic \[ax^2 + bx + c = 0\] multiply a and c in your question 2 x -15 = -30 find the factors of -30 that add to -1...... -6 and 5 then set up the solution in this manner (ax + factor 1)(ax + factor2) (2x - 6)(2x +5) ------------------------- =0 -------------- = 0 a 2 remove any common factors from the binomials 2(x - 3)(2x + 5) -------------- =0 2 and wow there is a common factor that cancels... and leaves you with a factorised quadratic. and allow you to find the zeros.